# KnowledgeSmith: Uncovering Knowledge Updating in LLMs with Model Editing and Unlearning

### v1

@@ 5.1 Setup
Models. Our evaluation covers 66 families of LLMs with 1B to 123B parameters, leading to a total of $13$ models: LLaMA-3 (1B, 3B, 8B, 70B) (Meta, 2024), Qwen-3 (1.7B, 14B, 32B) (Team, 2025b), QwQ-32B (Team, 2025a), Mistral (24B, 123B) (Jiang et al., 2023), Gemma (2B, 7B) (Team, 2024), and DeepSeek-R1-0528-Qwen3-8B (DeepSeek-AI, 2025). This broad coverage enables us to study whether scaling behaviors and editing/unlearning performance generalize across architectures.
Implementation Details. We adopted AlphaEdit (Fang et al., 2025) and ReLearn (Xu et al., 2025). AlphaEdit is a state-of-the-art editor that has been shown to outperform prior methods such as MEMIT(Meng et al., 2022b) and ROME(Meng et al., 2022a) in editing tasks, while ReLearn represents a leading approach to unlearning. Importantly, our framework is method-agnostic and directly extensible to other baselines, making it straightforward to integrate additional methods. Unlike traditional unlearning approaches where the retain set corresponds to the original knowledge, in our redirection-based setup the retain set is defined as the post-updated knowledge, ensuring that the model preserves the rewritten fact rather than reverting to its prior belief. This redirection-based formulation aligns better with real-world scenarios where knowledge is updated rather than erased. Editing and unlearning were applied separately to leaf, intermediate, and root nodes of the knowledge graph, with training data sizes ranging from 11 to $10,000$ samples. This setup allowed us to systematically analyze the effect of both hierarchy depth and data scale on the success of editing and unlearning. For evaluation, since each knowledge probing question is multiple-choice, we report accuracy as the proportion of questions for which the model selects the correct choice. This metric directly reflects the model’s correctness in retrieving or updating the intended knowledge.
@@ 5.2 Comparative Analysis of Editing and Unlearning
Human learners expect hierarchical consistency: updating a root concept should cascade to its descendants, while modifying a leaf should remain localized. We evaluate this in LLMs by applying editing or unlearning at three hierarchy levels (root, intermediate, leaf) and measuring performance on both targeted and structurally related nodes. We quantify these effects using direct vs. multi-hop accuracy (Figure˜2) as a proxy for propagation metrics: the Collateral Change Ratio (CCR) captures over-spreading for editing, and the Residual Retention ($R$) captures under-spreading for unlearning (For the complete definitions of CCR and $R$, see Appendix˜B).
@@ 5.5 Failure Mode and Stress Testing
Existing studies describe errors such as incomplete forgetting or knowledge pollution in a fragmented way, without systematically characterizing the underlying mechanisms. Through our experiments on open-ended question answering, we observed that models fail for different reasons under editing and unlearning interventions. To capture these patterns, we propose a Unified Failure Mode Taxonomy that organizes observed errors into six categories (examples of each type in Section˜D.2): under-forgetting ($R$), over-spreading (CCR), conflict emergence (contradictions between updated and related knowledge), knowledge drift (performance degradation on unrelated tasks), instruction-following drop (reduced ability to follow complex instructions), and hallucination increase.
@@ 5.6 Theoretical Analysis
Unlearning as anisotropic scaling. By contrast, Figure˜7(b) shows that unlearning produces sharper downscaling of singular values, with less stable alignment of $U,V$ across layers. This indicates suppression of capacity in certain subspaces rather than a simple rotation. Thus, unlearning resembles an attenuation operator: it removes the ability to encode certain directions but does not reliably rotate them into new ones. This mechanism aligns with the observed under-spreading behavior (high $R$ in Section˜5.2), where forgetting remains localized and fails to propagate fully across related nodes.
@@ Appendix B Propagation Asymmetry Metrics and Algorithm
Residual Retention ($R$) $=\frac{1}{|\mathcal{Q}_{\text{related}}|}\sum_{x\in\mathcal{Q}_{\text{related}}}\mathbf{1}\big[\hat{y}_{\theta^{\prime}}(x)=y_{\theta}(x)\big],$ (5)
@@ D.2 Failure Mode Examples
Biology (DNA) Under-forgetting ($R$) DNA year remains 1953 after update to 1955
@@
Economics (Phillips curve) Under-forgetting ($R$) Phillips curve still inflation-unemployment after update
@@
History (Declaration) Under-forgetting ($R$) Declaration year still 1776 after update to 1777
@@
Physics (General Relativity) Under-forgetting ($R$) GR year remains 1915 after update to 1920
@@ Appendix F Model Similarity Result
Computationally Efficiency. For the same model on a target dataset of $10,000$ examples, unlearning typically completes in about $1.5$ hours on an NVIDIA H100. Knowledge editing is more resource-intensive (roughly 66 hours). This additional cost highlights the heavier computational demands of precise factual editing.

### v2

@@ 5.1 Setup
Models. Our evaluation covers $6$ families of LLMs with 1B to 123B parameters, leading to a total of $13$ models: LLaMA-3 (1B, 3B, 8B, 70B) (Meta, 2024), Qwen-3 (1.7B, 14B, 32B) (Team, 2025b), QwQ-32B (Team, 2025a), Mistral (24B, 123B) (Jiang et al., 2023), Gemma (2B, 7B) (Team, 2024), and DeepSeek-R1-0528-Qwen3-8B (DeepSeek-AI, 2025). This broad coverage enables us to study whether scaling behaviors and editing/unlearning performance generalize across architectures.
Implementation Details. We adopted AlphaEdit (Fang et al., 2025) and ReLearn (Xu et al., 2025). AlphaEdit is a state-of-the-art editor that has been shown to outperform prior methods such as MEMIT(Meng et al., 2022b) and ROME(Meng et al., 2022a) in editing tasks, while ReLearn represents a leading approach to unlearning. Importantly, our framework is method-agnostic and directly extensible to other baselines, making it straightforward to integrate additional methods. Unlike traditional unlearning approaches where the retain set corresponds to the original knowledge, in our redirection-based setup the retain set is defined as the post-updated knowledge, ensuring that the model preserves the rewritten fact rather than reverting to its prior belief. This redirection-based formulation aligns better with real-world scenarios where knowledge is updated rather than erased. Editing and unlearning were applied separately to leaf, intermediate, and root nodes of the knowledge graph, with training data sizes ranging from $1$ to $10,000$ samples. This setup allowed us to systematically analyze the effect of both hierarchy depth and data scale on the success of editing and unlearning. For evaluation, since each knowledge probing question is multiple-choice, we report accuracy as the proportion of questions for which the model selects the correct choice. This metric directly reflects the model’s correctness in retrieving or updating the intended knowledge.
@@ 5.2 Comparative Analysis of Editing and Unlearning
Human learners expect hierarchical consistency: updating a root concept should cascade to its descendants, while modifying a leaf should remain localized. We evaluate this in LLMs by applying editing or unlearning at three hierarchy levels (root, intermediate, leaf) and measuring performance on both targeted and structurally related nodes. We quantify these effects using direct vs. multi-hop accuracy (Figure˜2) as a proxy for propagation metrics: the Collateral Change Ratio (CCR) captures over-spreading for editing, and the Residual Retention (RR) captures under-spreading for unlearning (For the complete definitions of CCR and RR, see Appendix˜B).
@@ 5.5 Failure Mode and Stress Testing
Existing studies describe errors such as incomplete forgetting or knowledge pollution in a fragmented way, without systematically characterizing the underlying mechanisms. Through our experiments on open-ended question answering, we observed that models fail for different reasons under editing and unlearning interventions. To capture these patterns, we propose a Unified Failure Mode Taxonomy that organizes observed errors into six categories (examples of each type in Section˜D.2): under-forgetting (RR), over-spreading (CCR), conflict emergence (contradictions between updated and related knowledge), knowledge drift (performance degradation on unrelated tasks), instruction-following drop (reduced ability to follow complex instructions), and hallucination increase.
@@ 5.6 Theoretical Analysis
Unlearning as anisotropic scaling. By contrast, Figure˜7(b) shows that unlearning produces sharper downscaling of singular values, with less stable alignment of $U,V$ across layers. This indicates suppression of capacity in certain subspaces rather than a simple rotation. Thus, unlearning resembles an attenuation operator: it removes the ability to encode certain directions but does not reliably rotate them into new ones. This mechanism aligns with the observed under-spreading behavior (high RR in Section˜5.2), where forgetting remains localized and fails to propagate fully across related nodes.
@@ Appendix B Propagation Asymmetry Metrics and Algorithm
Residual Retention (RR) $=\frac{1}{|\mathcal{Q}_{\text{related}}|}\sum_{x\in\mathcal{Q}_{\text{related}}}\mathbf{1}\big[\hat{y}_{\theta^{\prime}}(x)=y_{\theta}(x)\big],$ (5)
@@ D.2 Failure Mode Examples
Biology (DNA) Under-forgetting (RR) DNA year remains 1953 after update to 1955
@@
Economics (Phillips curve) Under-forgetting (RR) Phillips curve still inflation-unemployment after update
@@
History (Declaration) Under-forgetting (RR) Declaration year still 1776 after update to 1777
@@
Physics (General Relativity) Under-forgetting (RR) GR year remains 1915 after update to 1920
@@ Appendix F Model Similarity Result
Computationally Efficiency. For the same model on a target dataset of $10,000$ examples, unlearning typically completes in about $1.5$ hours on an NVIDIA H100. Knowledge editing is more resource-intensive (roughly $6$ hours). This additional cost highlights the heavier computational demands of precise factual editing.